OpenStudy (anonymous):

Solve x2 + 4x + 8 = 0

OpenStudy (anonymous):

You have to use the quadratic equation, do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

ax2+bx+c=0 Formula: (-b+-sqrt(b squared -4*a*c))/2a

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

a=1, b=4, c=8, plug 'em in

OpenStudy (anonymous):

x equals negative 2 plus or minus 4 I?

OpenStudy (anonymous):

BluFoot?

OpenStudy (anonymous):

+- stands for two possible resoults, both are correct

OpenStudy (anonymous):

\[ax^2 + bx + c = 0\]\[x^2 + 4x + 8 = 0\]a = 1 b = 4 c = 8 \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a } \] As BluFoot mentioned, just plug them in.\[x=\frac{ -(4) \pm \sqrt{(4)^2-4(1)(8)} }{ 2(1) } \]

OpenStudy (anonymous):

Now lets simplify that a bit. \[x=\frac{ -(4) \pm \sqrt{(4)^2-4(1)(8)} }{ 2(1)} \] \[x=\frac{ -(4) \pm \sqrt{16-32} }{ 2 }\]

OpenStudy (anonymous):

x equals negative 4 plus or minus 4 I?

OpenStudy (anonymous):

jazy?